### v1

# Learning Metrics that Maximise Power for Accelerated A/B-Tests

Abstract.
Online controlled experiments are a crucial tool to allow for confident decision-making in technology companies. A North Star metric is defined (such as long-term revenue or user retention), and system variants that statistically significantly improve on this metric in an A/B-test can be considered superior. North Star metrics are typically delayed and insensitive. As a result, the cost of experimentation is high: experiments need to run for a long time, and even then, type-II errors (i.e. false negatives) are prevalent.
We propose to tackle this by learning metrics from short-term signals that directly maximise the statistical power they harness with respect to the North Star. We show that existing approaches are prone to overfitting, in that higher average metric sensitivity does not imply improved type-II errors, and propose to instead minimise the $p$-values a metric would have produced on a log of past experiments. We collect such datasets from two social media applications with over 160 million Monthly Active Users each, totalling over 153 A/B-pairs. Empirical results show that we are able to increase statistical power by up to 78% when using our learnt metrics stand-alone, and by up to 210% when used in tandem with the North Star. Alternatively, we can obtain constant statistical power at a sample size that is down to 12% of what the North Star requires, significantly reducing the cost of experimentation.
1. Introduction & Motivation
Modern platforms on the web need to continuously make decisions about their product and user experience, which are often central to the business at hand. These decisions range from design and interface choices to back-end technology adoption and machine learning models that power personalisation. Online controlled experiments, the modern web-based extension of Randomised Controlled Trials (RCTs) (Rubin, 1974), provide an effective tool to allow for confident decision-making in this context (Kohavi et al., 2020) (bar some common pitfalls (Kohavi et al., 2022; Jeunen, 2023)).
A North Star metric is adopted, such as long-term revenue or user retention, and system variants that statistically significantly improve the North Star metric are considered superior to the tested alternative (Deng and Shi, 2016). Proper use of statistical hypothesis testing tools such as Welch’s $t$-test (WELCH, 1947), then allows us to define and measure statistical significance in a mathematically rigorous manner.
However effective this procedure is, it is far from efficient. Indeed, experiments typically need to run for a long time, and statistically significant changes to the North Star are scarce. This can either be due to false negatives (i.e. type-II error), or simply because the North Star is not moved by short-term experiments. In these cases, we need to resort to second-tier metrics (e.g. various types of user engagement signals) to make decisions instead. These problems are common in industry, as evidenced by a wide breadth of related work. A first line of research leverages control variates to reduce the variance of the North Star metric, directly reducing type-II errors by increasing sensitivity (Deng et al., 2013; Xie and Aurisset, 2016; Budylin et al., 2018; Poyarkov et al., 2016; Guo et al., 2021; Baweja et al., 2024). Another focuses on identifying second-tier “proxy” or “surrogate” metrics that are promising to consider instead of the North Star (Wang et al., 2022; Richardson et al., 2023; Tripuraneni et al., 2023), or to predict long-term effects from short-term data (Athey et al., 2019; Tang et al., 2022; Goffrier et al., 2023). Finally, several works learn metric combinations that maximise sensitivity (Deng and Shi, 2016; Kharitonov et al., 2017; Tripuraneni et al., 2023).
This paper synthesises, generalises and extends several of the aforementioned works into a general framework to learn A/B-testing metrics that maximise the statistical power they harness. We specifically extend the work of Kharitonov et al. (Kharitonov et al., 2017) to applications beyond web search, where the North Star can be delayed and insensitive. We highlight how their approach of maximising the average $z$-score does not accurately reflect downstream metric utility in our case, in that it does not penalise disagreement with the North Star sufficiently (i.e. type-III errors (Mosteller, 1948; Kaiser, 1960; Urbano et al., 2019)). Indeed: whilst this approach maximises the mean $z$-score, it does not necessarily improve the median $z$-score, and does not lead to improved statistical power in the form of reduced type-II error as a result.
Alternatively, optimising the learnt metric to minimise $p$-values —either directly or after applying a $\log$-transformation— more equitably divides gains over multiple experiments, leading to more statistically significant results instead of a few extremely significant results. Furthermore, we emphasise that learnt metrics are not meant to replace existing metrics, but rather to complement them. As such, their evaluation should be done through multiple hypothesis testing (with appropriate corrections (Shaffer, 1995)) if any of the North Star, available vetted proxies and surrogates, or learnt metrics are statistically significant under the considered treatment variant. We can then either adopt a conservative plug-in Bonferroni correction to temper type-I errors, or analyse synthetic A/A experiments to ensure the final procedure matches the required confidence level.
We empirically validate these insights through two dataset of past logged A/B results from large-scale short-video platforms with over 160 million monthly active users each: ShareChat and Moj. Experimental results highlight that our learnt metrics provide significant value to the business: learnt metrics can increase statistical power by up to 78% over the North Star, and up to 210% when used in tandem. Alternatively, if we wish to retain constant statistical power as we do under the North Star, we can do so with down to 12% of the original required sample size. This significantly reduces the cost of online experimentation to the business. Our learnt metrics are currently used for confident, high-velocity decision-making across ShareChat and Moj business units.
2. Background & Problem Setting
We deal with online controlled experiments, where two system variants $A$ and $B$ are deployed to a properly randomised sub-population of users, adhering to best practices (Kohavi et al., 2020; Jeunen, 2023).
For every system variant, for every experiment, we measure various metrics that describe how users interact with the platform. These metrics include types of implicit engagement (e.g. video-plays and watch time), as well as explicit engagement (e.g. likes and shares) as well as longer-term retention or revenue signals. For each metric, we log empirical means, variances and covariances (of the sample mean). For metrics $m_{i}$ with $1\leq i\leq N$, that is:
(1) $\bm{\mu}=[\mu_{1},\ldots\mu_{N}],\text{ and }\bm{\Sigma}=\begin{bmatrix}\sigma_{1}&\dots&\sigma_{1i}&\dots&\sigma_{1N}\\ \vdots&\ddots&\vdots&\ddots&\vdots\\ \sigma_{i1}&\dots&\sigma_{i}&\dots&\sigma_{iN}\\ \vdots&\ddots&\vdots&\ddots&\vdots\\ \sigma_{N1}&\dots&\sigma_{Ni}&\dots&\sigma_{N}\end{bmatrix}.$
Superscripts denote measurements pertaining to different variants in an experiment: e.g. $\bm{\mu}^{A}$ and $\bm{\mu}^{B}$.
2.1. Statistical Significance Testing
We want to assess whether the mean of metric $m_{i}$ is statistically significantly higher under variant $A$ compared to variant $B$. To this end, we define a confidence level $\alpha$ (often $\alpha\approx 0.05$), corresponding to the false-positive-rate we deem acceptable. Then, we apply Welch’s $t$-test. The test statistic (also known as the $z$-score) for metric $m_{i}$ and the given variants is given by:
(2) $z_{i}^{A\succ B}=\frac{\mu_{i}^{A}-\mu_{i}^{B}}{\sqrt{\sigma^{A}_{i}+\sigma^{B}_{i}}}.$
We then transform this to a $p$-value for a two-tailed test as:
(3) $p_{i}^{A\succ B}=2\cdot\min(\Phi(z_{i}^{A\succ B});1-\Phi(z_{i}^{A\succ B})).$
Here, $A\succ B$ denotes a partial ordering between variants, implying that $A$ is preferred over $B$. $\Phi(\cdot)$ represents the cumulative distribution function (CDF) for a standard Gaussian. For completeness, this CDF is given by:
(4) $\Phi(z)=\frac{1}{\sqrt{2\pi}}\int_{-\infty}^{z}e^{-\frac{t^{2}}{2}}{\rm d}t.$
When $p_{i}^{A\succ B}<\alpha$, we can confidently reject the null-hypothesis that $A$ and $B$ are equivalent w.r.t. the mean of metric $m_{i}$. Note that $z$-scores are signed, whereas two-tailed $p$-values are not. Indeed: $p_{i}^{A\succ B}\equiv p_{i}^{B\succ A}$, leaving room for faulty conclusions of directionality, known as type-III errors (Mosteller, 1948; Kaiser, 1960; Urbano et al., 2019). We discuss these later in depth.
A one-tailed $p$-value for the one-tailed null hypothesis $A\nsucc B$ is given by $1-\Phi(z_{i}^{A\succ B})$, and rejected when $<\frac{\alpha}{2}$. Throughout, we use two-tailed $p$-values unless mentioned otherwise.
2.1.1. $p$-value corrections
The above procedure is valid for a single metric, a single hypothesis, and importantly, a single decision. Nevertheless, this is not how experiments run in practice. Without explicit corrections on the $p$-values (or corresponding $z$-scores), violations of these assumptions lead to inflated false-positive-rates. We consider two common cases: a (conservative) multiple testing correction when an experiment has several treatments, and a sequential testing correction when experiments have no predetermined end-date or sample size at which to conclude. These corrections are applied as experiment-level corrections, to ensure that for any metric $m_{i}$ and variants $A,B$, the obtained $p$-values $p_{i}^{A\succ B}$ accurately reflect what they should reflect, yielding the specified coverage at varying confidence levels $\alpha$.
Multiple comparisons
Often, launched experiments will have multiple treatments deployed, leading to the infamous “multiple hypothesis testing” problem (Shaffer, 1995). We apply a Bonferroni correction to deal with this. When there are $T$ treatments, we consider a treatment to be statistically significantly different from control when a two-tailed $p<\frac{\alpha}{T}$, instead of the original $p<\alpha$ threshold.
We can equivalently apply this correction on $z$-scores instead, allowing us to directly compare $z$-scores across experiments with varying numbers of treatments. Recall that the percentile point function is the inverse of the CDF. We obtain a one-tailed $p$-value as $p=1-\Phi(z)$, and we reject the null hypothesis when $p<\frac{\alpha}{2}$. Now, instead, we reject when $p<\frac{\alpha}{2T}$. As such, computing corrected $z$-scores as $\bar{z}=z\frac{\Phi^{-1}(\frac{\alpha}{2})}{\Phi^{-1}(\frac{\alpha}{2T})}$ controls type-I errors effectively.
Always-Valid-Inference and peeking
A statistical test should only be performed once, at the end of an experiment. When the treatment effect is large, this implies we may have been able to conclude the experiment earlier. To this end, sequential hypothesis tests have been proposed in the literature (Wald, 1945). Modern versions make use of Always-Valid-Inference (AVI) (Howard et al., 2021) to allow for continuous peeking at intermediate results and making decisions based on them, whilst controlling type-I errors. Here, analogously, we can apply a correction on the $z$-scores as follows:
[…]
where $N_{AB}$ is the total number of samples over both variants combined. For a detailed motivation, see Schmit and Miller (Schmit and Miller, 2022).
These corrections are applied on a per-experiment level, both in the objective functions of methods introduced in the following Sections and when evaluating the metrics that they produce.
2.2. Learning Metrics that Maximise Sensitivity
The observation that we can learn parameters to maximise statistical sensitivity is not new. Yue et al. apply such ideas specifically for interleaving experiments in web search (Yue et al., 2010). Kharitonov et al. extend this to A/B-testing in web search, aiming to learn combinations of metrics that maximise the average $z$-score (Kharitonov et al., 2017). Deng and Shi discuss “lessons learned” from applying similar techniques (Deng and Shi, 2016). We introduce the approach presented by Kharitonov et al. (Kharitonov et al., 2017), as our proposed improvements build on their foundations.
We consider new metrics as linear transformations on $\bm{\mu}$:
(6) $\omega=\bm{\mu}w^{\intercal},\text{ where }w\in\mathbb{R}^{1\times N}.$
The advantage of restricting ourselves to linearity, it that we can write out the $z$-score of the new metric as a function of its weights:
(7) $z^{A\succ B}_{\omega}=\frac{\bm{\mu}^{A}w^{\intercal}-\bm{\mu}^{B}w^{\intercal}}{\sqrt{w\bm{\Sigma}^{A}w^{\intercal}+w\bm{\Sigma}^{B}w^{\intercal}}}.$
These $z$-scores can be used exactly as before to obtain $p$-values. An intuitive property of the $z$-score, is that a relative $z$-score of $r=\frac{z^{A\succ B}_{\omega}}{z^{A\succ B}_{i}}$ implies that $\omega$ requires a factor $r^{2}$ fewer samples to achieve the same confidence level as $m_{i}$ (Chapelle et al., 2012). This can directly be translated to the cost of experimentation, as it allows us to run experiments for shorter time-periods or on smaller sub-populations.
As such, it comes naturally to frame the objective as learning the weights $w$ that maximise the $z$-score on the training data. This training data consists of a set of experiments with pairs of variants =$\mathcal{E}=\{(A,B)_{i}\}_{i=1}^{|\mathcal{E}|}$. We consider three distinct relations between pairs of deployed system variants:
  1. (1)
    Known outcomes: $(A,B)\in\mathcal{E}^{+}$, where $A\succ B$,
  2. (2)
    Unknown outcomes: $(A,B)\in\mathcal{E}^{?}$, where $A~{}?~{}B$,
  3. (3)
    A/A outcomes: $(A,B)\in\mathcal{E}^{\simeq}$, where $A\simeq B$.
Here, $A\succ B$ implies that there is a known and vetted preference of variant $A$ over $B$ — typically because the North star or other guardrail metrics showed statistically significant improvements. We denote inconclusive experiments as $A~{}?~{}B$, implying statistically insignificant outcomes on the North Star. In rare cases, the North star might have gone up at the expense of important guardrail metrics, rendering conclusions ambiguous. Finally, $A\simeq B$ represents A/A experiments, where we know the null hypothesis to hold true (by design). The first set of experiments is used to measure type-III errors. Known and unknown outcomes are used to measure type-II errors, and A/A experiments can inform us about type-I errors. This dataset of past A/B experiments is collected and labelled by hand, from natural experiments occurring on the platform over time.
2.2.1. Optimising Metric Weights with a Geometric Heuristic
Note that $z$ as a function of $w$ is scale-free. That is, the direction of the weight vector $w$ matters, but its scale does not. As Kharitonov et al. write (Kharitonov et al., 2017), we can compute the optimal direction of $w$ using the Lagrange multipliers method, to obtain:
(8) $w^{\star}_{A\succ B}\propto(\bm{\Sigma}^{A}+\bm{\Sigma}^{B}+\epsilon I)^{-1}(\bm{\mu}^{A}-\bm{\mu}^{B}).$
Here, $\epsilon\in\mathbb{R}$ is a small number to ensure that the matrix to be inverted is not singular. Kharitonov et al. fix this value at $\epsilon=0.01$ and never adjust it throughout the paper. We wish to highlight that technique is known as Ledoit-Wolf shrinkage (Ledoit and Wolf, 2004, 2020), and that it can have substantial influence on the obtained direction. Indeed: it acts as a regularisation term pushing the weights closer to $w=(\bm{\mu}^{A}-\bm{\mu}^{B})$. This can be seen by observing that as ϵinfitalic-$\epsilon\to\inf$, the inverse becomes $\frac{1}{\epsilon}I$, and the solution hence becomes $w=\frac{(\bm{\mu}^{A}-\bm{\mu}^{B})}{\epsilon}$. As we only care about the direction we can ignore the denominator. To ensure fair comparison, we also set $\epsilon=0.01$. Exploring the effects of Ledoit-Wolf shrinkage as a regularisation technique where $\epsilon$ is a hyper-parameter, gives an interesting avenue for future work.
In order to include observations from multiple experiments into a single set of learned weights, they propose to compute the optimal direction per experiment, normalise, and average the weights:
(9) $w^{\star}_{\mathcal{E}^{+}}=\frac{1}{|\mathcal{E}^{+}|}\sum_{(A,B)\in\mathcal{E}^{+}}\frac{w^{\star}_{A\succ B}}{\left\|w^{\star}_{A\succ B}\right\|_{2}}.$
Whilst this procedure provides no guarantees about the sensitivity of the obtained metric on the overall set of experiments, it is efficient to compute and provides a strong baseline method.
2.2.2. Optimising Metric Weights via Gradient Ascent on $z$-scores
A more principled approach is to cast the above as an optimisation problem. The objective function for this optimisation problem consists of three parts. First, we wish to maximise the $z$-score for all variant pairs with known outcomes:
(10) $\begin{split}\mathcal{L}_{z}^{+}(w;\mathcal{E}^{+})=\frac{1}{|\mathcal{E}^{+}|}\sum_{(A,B)\in\mathcal{E}^{+}}z^{A\succ B}_{\omega}.\end{split}$
Second, we wish to maximise the absolute $z$-score for all variant pairs with inconclusive outcomes under the North Star:
(11) $\mathcal{L}_{z}^{?}(w;\mathcal{E}^{?})=\frac{1}{|\mathcal{E}^{?}|}\sum_{(A,B)\in\mathcal{E}^{?}}\left|z^{A\succ B}_{\omega}\right|.$
Third, we wish to minimise the absolute $z$-score for all variant pairs that are equivalent (i.e. A/A-pairs):
(12) $\mathcal{L}_{z}^{\simeq}(w;\mathcal{E}^{\simeq})=\frac{1}{|\mathcal{E}^{\simeq}|}\sum_{(A,B)\in\mathcal{E}^{\simeq}}\left|z^{A\succ B}_{\omega}\right|.$
This gives rise to a combined objective as a weighted average:
(13) $\mathcal{L}_{z}(w;\mathcal{E})=\mathcal{L}_{z}^{+}(w;\mathcal{E}^{+})+\lambda^{?}\mathcal{L}_{z}^{?}(w;\mathcal{E}^{?})-\lambda^{\simeq}\mathcal{L}_{z}^{\simeq}(w;\mathcal{E}^{\simeq}).$
Kharitonov et al. demonstrate that, for a variety of different metrics in a web search engine, these approaches can exhibit improved sensitivity (Kharitonov et al., 2017). We apply this method to learn instantaneously available proxies to a delayed North Star metric in general scenarios, and propose several extensions, detailed in the following Sections.
3. Methodology & Contributions
3.1. Learning Metrics that Maximise Power
When directly optimising $z$-scores, an implicit assumption is made that the utility we obtain from increased $z$-scores is linear. This is seldom a truthful characterisation of reality, considering how we wish to actually use these metrics downstream. We provide a toy example in Table 1, reporting $z$-scores and one-tailed $p$-values for two experiments and three possible metrics $m_{1},m_{2},m_{3}$, inspired by real data. In this toy example, we know that $A\succ B$, based on a hypothetical North Star metric. Nevertheless, as we do not know this beforehand, we typically test for the null hypothesis $A\simeq B$ with two-tailed $p$-values. In the table, this means that the outcome is statistically significant if the reported one-tailed values are $p<\frac{\alpha}{2}$ or $p>1-\frac{\alpha}{2}$. We report the power of every metric at varying confidence levels $\alpha\in\{0.05,0.01\}$, reporting whether (i) the null hypothesis is correctly rejected ($p<\frac{\alpha}{2}$, ), (ii) the outcome is inconclusive (?), i.e. a type-II error, or (iii) the null hypothesis is rejected, but for the wrong reason ($p>1-\frac{\alpha}{2}$, $\times$).
This latter case is deeply problematic, as it signifies disagreement with the North Star. Such errors have been described as type-III errors in the statistical literature (Mosteller, 1948; Kaiser, 1960; Urbano et al., 2019). Naturally, we would rather have a metric that fails to reject the null than one that confidently declares a faulty variant to be superior. Indeed, Deng and Shi argue that both directionality and sensitivity are desirable attributes for any metric (Deng and Shi, 2016). Nevertheless, considering candidate metrics in Table 1, type-III errors are not sufficiently penalised by the average $z$-score: metric $m_{3}$ maximises this objective despite yielding statistical power that is on par with a coin flip.
Directly maximising power might prove cumbersome, as it is essentially a discrete step-function w.r.t. the $z$-score, dependent on the confidence level $\alpha$. Instead, it comes natural to minimise the one-tailed $p$-value reported in Table 1. Indeed, the $p$-value transformation models diminishing returns for high $z$-scores, which allows type-III errors to be sufficiently penalised. When considering this objective, $m_{3}$ is clearly suboptimal whilst $m_{2}$ is preferred.
Note that the change in objective would not affect the geometric heuristic as described in Section 2.2.1. As we simply apply a monotonic transformation on the $z$-scores, the weight direction that maximises the $z$-score equivalently minimises its $p$-value. When learning via gradient descent, however, the $p$-value transformation affects how we aggregate and attribute gains over different input samples. This allows us to stop focusing on increasing sensitivity for experiments that are already “sensitive enough”, and more equitably consider all experiments in the training data.
$z$-score $p$-value Power
$\alpha=0.05$ $\alpha=0.01$
[…]
Table 1. An example setting highlighting that maximising average $z$-scores ($m_{1}$) does not imply improved statistical power. Statistical power is dependent on the significance level $\alpha$, $p$-values provide a good proxy to optimise for. We observe similar trade-offs from real past A/B experiments.
This change in objective provides an intuitive and efficient extension to existing approaches, allowing us to directly optimise the confidence we have to correctly reject the null-hypothesis. For known outcomes, the loss function is given by:
(14) $\begin{split}\mathcal{L}_{p}^{+}(w;\mathcal{E}^{+})=\frac{1}{|\mathcal{E}^{+}|}\sum_{(A,B)\in\mathcal{E}^{+}}1-\Phi(z^{A\succ B}_{\omega})\\ =\frac{1}{|\mathcal{E}^{+}|}\sum_{(A,B)\in\mathcal{E}^{+}}1-\Phi\left(\frac{\bm{\mu}^{A}w^{\intercal}-\bm{\mu}^{B}w^{\intercal}}{\sqrt{w\bm{\Sigma}^{A}w^{\intercal}+w\bm{\Sigma}^{B}w^{\intercal}}}\right),\end{split}$
and analogously extended to unknown $\mathcal{L}_{p}^{?}$ and A/A-outcomes $\mathcal{L}_{p}^{\simeq}$. Nevertheless, we wish to point out that we only want to maximise $p$-values for A/A-outcomes if type-I error becomes problematic. As we will show empirically in Section 4.3, this is not a problem we encounter. For this reason, we set $\lambda^{\simeq}\equiv 0$.
Note that whilst direct optimisation of $p$-values is an improvement over myopic consideration of $z$-scores, there is another caveat: the “worst-case” loss of a type-III error is bounded at 1, which does not reflect our true utility function: metrics that disagree with the North Star are far less reliable than those that simply remain inconclusive. As such, we also consider another variant of the objective, where $\bar{p}=-p\log(1-p)$. Figure 1 provides visual intuition to clarify how this monotonic transformation on the $p$-values more heavily penalises type-III errors, whilst retaining the optimum. From a theoretical perspective, this function provides a convex relaxation for minimising the number of type-III errors a metric produces. As a result, we expect this surrogate to exhibit strong generalisation. We refer to this objective as minimising the $\log p$-value.
Figure 1. Visualising our proposed optimisation objectives for learnt metrics, as a function of their $z$-score.
Note that one could envision further extensions here where the confidence level $\alpha$ is directly incorporated into the objective function to maximise statistical power at a given confidence level $\alpha$. Nevertheless, we conjecture that their discrete nature might hamper effective optimisation and generalisation, compared to the strictly convex and smooth surrogate we obtain from the $\log p$-value.
3.2. Accelerated Convergence for Scale-Free Objectives via Spherical Regularisation
The objective functions we describe —either $z$-scores or $(\log)p$-values— are scale-free w.r.t. the weights that are being optimised. As a result, out-of-the-box gradient-based optimisation techniques are not well-equipped to handle this efficiently.
Consider a simple toy example where we have two observed metrics for an experiment with a known preference $A\succ B$, and:
(15) $\bm{\mu}^{A}=[1.0,1.0],\qquad\bm{\mu}^{B}=[0.5,0.5],\qquad\bm{\Sigma}^{A}=\bm{\Sigma}^{B}=I.$
For this low-dimensional problem, we can visualise the $z$-score as a function of the metric weights in a contour plot, as shown in Figure 1(a).
(a) Original scale-free objective function.
[…]
Figure 2. Directly maximising $z$-scores yields a scale-free objective that is not amenable to efficient optimisation with gradient ascent (left). Spherical regularisation retains all optima, whilst providing a gradient direction that is more aligned (right).
Here, it becomes visually clear that whilst the direction of the $w=[w_{1},w_{2}]$ vector matters, its scale does not. The consequence is that the gradient vectors w.r.t. the objective on the right-hand plot can lead to slow convergence, even in this concave objective. Indeed, for poor initialisation in the bottom left quadrant (e.g. $w=[-1,-2]$), the gradient direction is perpendicular w.r.t. the optima.
Recent work makes a similar observation for discrete scale-free objectives as they appear in ranking problems (Ustimenko and Prokhorenkova, 2020). They propose to adopt projected gradient descent, normalising the gradients before every update. Whilst effective, in our setting we would prefer to use out-of-the-box optimisation methods for practitioners’ ease-of-use. Instead, we introduce a simple regularisation term that represents the distance between the scale of the $w$ vector and a hyper-sphere:
(16) $\mathcal{L}_{\lVert w\rVert}=-\delta\left(N-\lVert w\rVert_{2}^{2}\right)^{2}.$
All optima for this objective function are also optima to the original function—but the gradient field is more amenable to iterative gradient-based optimisation techniques. Figure 1(b) visualises how this transforms the loss surface. Under this regularised objective, it is visually clear that gradient-based optimisation methods are likely to exhibit faster convergence. Our empirical results confirm this, for a variety of initialisation weights and learning objectives.
4. Experiments & Discussion
To empirically validate the methods proposed in this work, we require a dataset containing logged metrics (sample means, their variances and covariances), together with preference orderings over competing system variants that were collected from real-world A/B-tests, ideally spanning large user-bases and several weeks.
Existing work on this topic used a private dataset from Yandex focused on web search experiments that ran between 2013–2014 (Kharitonov et al., 2017). They report type-I and -II errors for 8 metrics and a fixed 5% confidence level, over 118 A/B-tests and 472 A/A-tests.
In this work, we consider more general metrics that are relevant for use-cases beyond web search (i.a. user retention and various engagement signals). Furthermore, we report type-I/II/III errors at varying confidence levels, providing insights into the learnt metrics’ behaviour. For this, we leverage logs of past A/B-experiments on two large-scale short-video platforms with over 160 million monthly active users each: ShareChat and Moj. The datasets consist of 153 A/B-experiments total that ran in 2023, and over 25 000 A/A-pairs. In total, we have access to roughly 100 metrics detailing various types of interactions with the platform, engagements, and delayed retention signals. Because our dataset is limited in size (a natural consequence of the problem domain), we are bound to overfit when using all available metrics as input features. As such, we limit ourselves to 10 input metrics to learn from, and evaluate them w.r.t. the delayed North Star. This feature selection step also ensures that our linear model consists of fewer parameters, which increases practitioners’ and business stakeholders’ trust in its output. We focus on non-delayed signals, including activity metrics such as the number of sessions and active days, and counters for positive and negative feedback engagements of various types. These are selected through an analysis of their type-I/II/III errors w.r.t. the North Star, as well as their $z$-scores: focusing on metrics with high sensitivity and limited disagreement. The research questions we wish to answer empirically using this data, are the following:
RQ1:
Do learnt metrics effectively improve on their objectives?
RQ2:
How do learnt metrics behave in terms of type-III errors?
RQ3:
How do learnt metrics’ type-I/II errors behave when considered as stand-alone evaluation metrics?
RQ4:
How do learnt metrics’ type-I/II errors behave when used in conjunction with the North Star and top proxy metrics?
RQ5:
How do learnt metrics influence required sample sizes when used in conjunction with the North Star and top proxy metrics?
RQ6:
Do we observe accelerated convergence over varying objectives via the proposed spherical regularisation technique?
We report results for the ShareChat platform in what follows, and provide further empirical results for Moj in Appendix A.
4.1. Effectiveness of Learnt Metrics (RQ1)
We learn and evaluate metrics through leave-one-out cross-validation: for every experiment, we train a model on all other experiments and evaluate the $z$-score (Eq. 7) and $p$-value (Eq. 3) the metric yields for the held-out experiment. We report the mean and median $z$-scores and $p$-values we obtain for all A/B-pairs with known outcomes (i.e. $\mathcal{E}^{+}$) in Table 2. Best performers for every column (either maximising $z$-scores or minimising $p$-values) are highlighted in boldface. Empirical observations match our theoretical expectations: whilst the $z$-score objective does effectively maximise the average $z$-score, it is the worst performer for both mean and median $p$-values, and even the median $z$-score. Our proposed $\log p$-value objective effectively improves both the median $z$-score and $p$-value over alternatives.
[…]
Mean Median Mean Median
[…]
Table 2. Sensitivity results for learnt metrics, computed via leave-one-out cross-validation on all experiments with known outcomes. We observe that minimising the $\log p$-value effectively improves median sensitivity over alternatives.
Figure 3. Learnt metrics and their agreement with known North Star outcomes over varying confidence levels. A learnt metric that maximises the average $z$-score exhibits significant type-III error, which can be alleviated by minimising $p$-values instead.
(a) False-positive rate (type-I error) and false-negative rate (type-II error, $1-$ power) for stand-alone metrics.
(b) False-positive rate (type-I error) and false-negative rate (type-II error, $1-$ power) for sets of metrics, after Bonferroni correction.
Figure 4. When considering only learnt metrics, we improve type-II error significantly without hurting specificity. At $\alpha=0.05$, we increase statistical power by 67% (upper plot). In conjunction with the North Star and top proxy metrics, Bonferroni corrections are slightly conservative (type-I error $<\alpha$), and allow us to improve statistical power by 133% for $\alpha=0.05$ (lower plot).
4.2. Agreement with the North Star (RQ2)
From the obtained $z$-scores and $p$-values summarised in Table 2, we can additionally derive (dis-)agreement with the North Star, for varying confidence levels $\alpha$. We visualise these results in Figure 3: if the obtained $p$-value under a learnt metric is lower than $\alpha$, that metric yields a statistically significant result (agreement). If the obtained $p$-value for the alternative hypothesis (i.e. $B\succ A$ when we know $A\succ B$) is lower than $\alpha$, we have statistically significant disagreement, or type-III error. This is a capital sin we wish to avoid at all costs, as it severely diminishes the trust we can put in the learnt metric. If the $p$-value reveals a statistically insignificant result, we say the result is inconclusive, implying a type-II error. We observe that both the $z$-score-maximising metric and the heuristic approach fail to steer clear from type-III error. Optimising ($\log$)$p$-values instead, alleviates this issue. For this reason, we only consider these metrics for further evaluation. Indeed: an analysis of type-II error is rendered meaningless when type-III errors are present.
Figure 6(a) in Appendix A highlights that for the Moj platform as well, type-III errors are common in the case of $z$-score-maximising or heuristic metrics. As such, we only consider ($\log$)$p$-values to assess increases in statistical power and potential reductions to the cost of running online experiments.
4.3. Power Increase from Learnt Metrics (RQ3–4)
Until now, we have leveraged experiments with known outcomes to assess sensitivity and agreement with the North Star. Now, we additionally consider A/A-experiments ($\mathcal{E}^{\simeq}$) and experiments with unknown outcomes ($\mathcal{E}^{?}$) to measure type-I and type-II error respectively. We measure this for the North Star, for the best-performing proxy metric that serves as input to the learnt metrics, and for learnt metrics that exhibit no empirical disagreement with the North Star. We plot the type-I error (i.e. fraction of A/A-pairs in $\mathcal{E}^{\simeq}$ that are statistically significant at confidence level $\alpha$) and the type-II error (i.e. fraction of A/B-pairs in {$\{\mathcal{E}^{+}\cup\mathcal{E}^{?}\}$ that are statistically insignificant at confidence level $\alpha$) for varying values of $\alpha$ in Figure 3(a). We observe that we are able to significantly reduce type-II errors compared to the North Star (up to 78%), whilst keeping type-I errors at the required level (i.e. $\alpha$). However, we also observe that the type-II error we obtain when using learnt metrics does not significantly improve over the top proxy metric, when considered in isolation.
Nonetheless, this is not how evaluation metrics are used in practice. Indeed, we track several metrics and can draw conclusions if any of them are statistically significant. As such, metrics should be evaluated on their complementary sensitivity. That is, we compute $p$-values for a set of metrics, apply a Bonferroni correction, and assess statistical significance. The statistical power that we obtain through this procedure is visualised in Figure 3(b). We consider either the North Star in isolation, the North Star in conjunction with the top-proxy, or a further combination with any learnt metric. Here, we observe that the learnt metric provides a substantial increase in statistical power: statistical power (i.e. $1-$ type-II error) is increased by up to 210% compared to the North Star alone, and 25–30% over the North Star plus proxies. Furthermore, as the Bonferroni correction is slightly conservative, we observe lower than expected type-I error for higher confidence levels $\alpha$. This implies that a more fine-grained multiple testing correction can further improve statistical power. We empirically observe that this works as expected, but its effects are negligible in practice.
4.4. Cost Reduction from Learnt Metrics (RQ5)
So far, we have shown that metrics learnt to minimise ($\log$)$p$-values are effective at improving sensitivity (Table 2), whilst minimising type-III error (Figure 3) and improving statistical power (Figure 4).
On one hand, powerful learnt metrics can lead to more confident decisions from statistically significant A/B-test outcomes. Another view is that we could make the same amount of decisions based on fewer data points, as we reach statistical significance with smaller sample sizes. This implies a cost reduction, as we can run experiments either on smaller portions of user traffic or for shorter periods of time, directly impacting experimentation velocity.
This reduction in required sample size is equal to the square of the relative $z$-score (Chapelle et al., 2012; Kharitonov et al., 2017). We visualise this quantity in Figure 5, for varying confidence levels $\alpha$, for the same Bonferroni-corrected procedure as Figure 3(b). To obtain a $z$-score for a set of metrics, we simply take the maximal score and apply a Bonferroni correction to it as laid out in Section 2.1.1. Note that this procedure depends on $\alpha$, explaining the slope in Figure 5. We observe that our learnt metrics can achieve the same level of statistical confidence as the North Star with up to 8 times fewer samples, i.e. a reduction down to $12.5\%$. This significantly reduces the cost of experimentation for the business, further strengthening the case for our learnt metrics.
Figure 5. When considering learnt metrics in conjunction with the North Star and top proxy metrics, we require significantly reduced sample sizes to obtain the same statistical confidence level as we would get from the North Star.
4.5. Spherical Regularisation (RQ6)
[…]
Figure 6 visualises the evolution of the learning objective over optimisation steps, for all mentioned learning objectives, initialisation strategies, and regularisation strengths. We observe that the method is robust, significantly improving convergence speed for all settings, requiring up to 40% fewer iterations until convergence is reached. This positively influences the practical utility of the learnt metric pipeline for researchers and practitioners.
Figure 6. Spherical regularisation significantly accelerates convergence for all considered objectives, up to 40%.
5. Insights from Learnt Metrics
In this Section, we briefly discuss insights that arose through our empirical evaluation of all metrics: the North Star, classical surrogates and proxies, as well as learnt metrics. These insights are specific to our platforms, but we believe they can contribute to a general intuition and understanding of metrics for online content platforms and broader application areas.
Ratio metrics are easily fooled.
Often, important metrics can be framed as a ratio of the means (or sums) of two existing metrics (Baweja et al., 2024; Budylin et al., 2018). Examples include click-through rate (i.e. clicks / impressions), variants of user retention (i.e. retained users / active users), or general engagement ratios (e.g. likes / video-plays). We observe that, whilst these metrics can be important from a business perspective, they typically exhibit significant type-III errors w.r.t. the North Star. Indeed, in the examples above both the numerator and denominator represent positive signals we wish to increase. Suppose an online experiment increases the number of video-plays by $Y\%$, and the overall number of likes by $X\%<Y\%$. These two positive signals will lead to a decreasing ratio, whilst we are likely to still prefer the treatment w.r.t. the North Star if $X$ and $Y$ are substantially large.
We believe this is connected to common offline ranking evaluation metrics prevalent in the recommender systems field (Steck, 2013; Jeunen, 2019). Indeed, such metrics are cumulative in nature, optimising overall value instead of some notion of value-per-item (Jeunen et al., 2023).
User-level aggregations conquer general counters.
In the previous example, we describe general count metrics for the number of likes and the number of video-play events. User behaviour on online platforms often follows a power-law distribution: a few “power users” generate the majority of such events (Chi, 2020). As a result, such metrics are easily skewed, and they are not guaranteed to accurately reflect improvements for the full population of users. Aggregating such counters per users (e.g. count the number of days a user has at least $X$ video-plays) instead of using raw event counters, provides strong and sensitive proxies to the North Star.
Interestingly, this framing is reminiscent of recall, as we effectively measure the coverage of users about whom we have positive signals. Recall metrics are again strongly connected to offline evaluation practices in recommender systems, especially in the first stage of two-stage systems common in industry (Covington et al., 2016; Ma et al., 2020).
6. Conclusions & Outlook
A/B-testing is a crucial tool for decision-making in online businesses, and it has widely been adopted as a go-to approach to allow for continuous system improvements. Notwithstanding their popularity, online experiments are often expensive to perform. Indeed: many experiments lead to statistically insignificant outcomes, presenting an obstacle for confident decision-making. Experiments that do lead to significant outcomes are costly too: by their very definition, a portion of user traffic interacts with a sub-optimal system variant. As such, we want to maximise the number of decisions we can make based on the experiments we run, and we want to minimise the required sample size for statistically significant outcomes. In this work, we achieve this by learning metrics that maximise the statistical power they harness. We present novel learning objectives for such metrics, and provide a thorough evaluation of the effectiveness of our proposed approaches. Our learnt metrics are currently used for confident, high-velocity decision-making across ShareChat and Moj business units.
Appendix A Additional Experimental Results
To further empirically validate our theoretical insights w.r.t. the proposed methods, we repeat the experiments reported in Section 4 on data collected for the Moj platform, and reproduce Figures 35. Results are visualised in Figure 7 Observations match our expectations, further strengthening trust in the replicability of our results.
All improvements in sensitivity and statistical power are a similar order of magnitude as those for ShareChat: learnt metrics that minimise ($\log$)$p$-values can substantially reduce type-II/III errors without affecting type-I errors. We observe an improvement over ShareChat data in Figure 6(d): learnt metrics for Moj exhibit a 12-fold reduction in the sample size that is required to attain constant statistical confidence as to the North Star.